The stockholders’ equity of Crater, Inc. at December 31, 2017, appears below:

Common stock, $25 par value, 200,000 shares authorized,135,000 shares issued and outstanding

3,375,000

Paid-in capital in excess of par value

1,750,000

Retained earnings

2,575,000

During 2018, the following transactions occurred:

Jan 1                 Sold 50,000 Shares of preferred stock at $75, par is $50, stated rate of return is 5%

May 10:            Declared a 10% common stock dividend when market value was $50 per share

June 15              Repurchased 5,000 share of common stock at $32

August 31:        Issued the stock dividend declared on May 10.

December 10: Declared a cash dividend of $300,000 to be paid in December 2018.

December 31: Recognized net income of $ 1,630,500.

  1. What is the amount, if any of the APIC for the preferred stock? ____________

  1. What is the amount is the treasury stock, if any? ____________

  1. What is the dividend per share of common stock, if any? ___________

  1. How many shares of common stock are outstanding at year end? _______

  1. What is the ending balance of Retained earnings? ____________

  1. What is the ending balance of total equity? ______________
Answer #1

a) APIC for the preferred stock = (75-50)*50000 = $1250000

b) Amount is the treasury stock = 5000*32 = 160000

c) Dividend per share of common stock =

Preferred dividend = 50000*50*5% = 125000

Common Stock = 300000-125000 = 175000

Share outstanding = 135000*1.1-5000 = 143500

Dividend per share of common Stock = 175000/143500 = 1.22 per share

d) Common stock outstanding = 135000*1.1-5000 = 143500

e) Ending balance of retained earnings = 2575000-(13500*50)-300000+1630500 = 3230500

f) Total equity = 3375000+1750000+(50000*75)+3230500+(13500*50)-(5000*32) = 41420500